OpenStudy (moongazer):

find the value of theta Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (moongazer):

\[3\cos ^{2}\theta-7\cos \theta+4=0\]

OpenStudy (anonymous):

Let cos(x) = b 3b^2 - 7b + 4 = 0 Solve b

OpenStudy (anonymous):

WTF wikipedia?

OpenStudy (anonymous):

You should probably get four answers

OpenStudy (moongazer):

b=1, b=4/3

OpenStudy (moongazer):

\[0\le \theta \le 2 \pi\]

OpenStudy (moongazer):

could there be an undefined solution for the b=4/3

OpenStudy (anonymous):

cos theta = 1 theta = 0, 2pi

OpenStudy (moongazer):

because there is no cosine that has a value of 4/3 right?

OpenStudy (moongazer):

so that is the only answer?

OpenStudy (anonymous):

It's undefined because cos theta = Adjacent over hypotenuse hypotenuse is never greater than the adjacent (leg)

OpenStudy (anonymous):

probably Remember to press good answer.

OpenStudy (moongazer):

so cos, sin and tan should never have a value greater than 1?

OpenStudy (anonymous):

For tan you can But not for cos or sin

OpenStudy (moongazer):

ohh i see because tan =y/x Thanks!